What operation does "@" do?
Study these examples
a. 4@5=3 b. 7@2=12 c.5@2=8 d. 3@6=0
e. 3@2=4 f. 5@6=4 c.2@3=q d 5@5=5
Solve the following
1.) 1@0=___
2.) 4@2=___
3.) 7@___=7
4. ___@9=7
Please answer the blank spaces or ___

Answer by solve_for_x(190)

The @ operation multiplies the first number by 2, and then subtracts the second number.

4@5 = 4*2 - 5 = 8 - 5 = 3

7@2 = 7*2 - 2 = 14 - 2 = 12

5@2 = 5*2 - 2 = 10 - 2 = 8

The remaining problems are then:

1. 1@0 = 1*2 - 0 = 2 - 0 = 2

2. 4@2 = 4*2 - 2 = 8 - 2 = 6

3. 7@___ = 7 --> 7*2 - ___ = 7 --> 14 - ___ = 7

4. ___@ 9 = 7 --> ___*2 - 9 = 7 --> ___*2 = 16

I'll leave it to you to determine the values for the blank spaces in #3 and #4.